OpenStudy (highschoolmom2010):

Find the solutions of the equation. 1/2x^2 - x + 5 = 0

OpenStudy (anonymous):

\[1/2x^2 - x + 5 = 0\] easier to multiply by two and solve \[x^2-2x+10=0\]

OpenStudy (anonymous):

i don't think this one factors, so you can either complete the square (easiest in this case because 2 is even) or you use the quadratic formula \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=-2,c=10\]